Express V= 2t² + 5t + -5 as a linear combination of the polynomials
P1= t² + 2t + 1
P2= 2t + 5t + 4
P3 = t² + 3t + 6
1
Expert's answer
2020-11-11T08:39:39-0500

We assume that "P_2=2t^2+5t+4". "V" can be expressed in the following way: "V=c_1P_1+c_2P_2+c_3P_3" , "c_1,c_2,c_3\in \mathbb{C}" . Coefficients satisfy the following system:

"\left\{\begin{matrix}
 c_1+2c_2+c_3=2 \\
 2c_1+5c_2+3c_3=5 \\
c_1+4c_2+6c_3=-5
\end{matrix}\right."

We multiply the first equation by "(-2)" and add to the second one. Then, we multiply the first equation by "(-1)" and add to the third one. As a result, we obtain we eliminate "c_1" from all equations and obtain the system for "c_2,c_3:"

"\left\{\begin{matrix}
c_2+c_3=1 \\
2c_2+5c_3=-7
\end{matrix}\right."

We multiply the first equation by "(-2)" and add to the second one. We obtain that "c_3=-3" . Then, from the first equation of the latter system we get: "c_2=4". From the first system we get "c_1=-3" . Thus, the solution is: "c_1=-3,c_2=4,c_3=-3" . The answer is: "V=-3P_1+4P_2-3P_3" .